Given that P is a point on the circle x^2 +y^2 - 4x +2y - 4=0. Find the longest distance from P to the y-axis.

1
Expert's answer
2021-11-02T19:04:00-0400

(x-a) 2 + (y-b) 2=r2

(a, b) is the center of the circle

X2+y2-4x+2y-4=0

X2-4x+(-4/2)2+y2+2y+(2/2)2=4+(-4/2)2+(2/2)2

(x-2)2+(y+1)2=9

-a=-2

a=2

-b=1

b=-1

r2=9

r=3

The longest Distance =3+2=5

Ans=5units